# On the Role of Context in Reading Time Prediction

### v3

@@ 2.2 Frequency as a Predictor of Reading Time
where the normalizing constant ${{Z}_{{q_{\mathrm{H}}}}}$ is necessarily finite for language models of finite expected length. [^3] As a counterexample, consider an LM ${{p_{\mathrm{H}}}}$ with ${\Sigma}=\{a\}$ and ${{p_{\mathrm{H}}}}(a^{n})=\frac{1}{\pi^{2}/6}\cdot\frac{1}{n^{2}}$ for $n\in\mathbb{Z}_{>0}$, i.e., where the probabilities are globally normalized by $\frac{\pi^{2}}{6}$, the solution to the Basel problem. The expected count of $a$ would depend on $\sum_{n=1}^{\infty}\frac{1}{n^{2}}\cdot n$, which is divergent. Thus, ${{Z}_{{q_{\mathrm{H}}}}}=\infty$. Then, given an LM ${{p_{\mathrm{H}}}}$ with a unigram LM ${{q_{\mathrm{H}}}}$, the unigram surprisal is given by
@@ 4 Variance Explained by Context
We fit three linear models using ordinary least squares, with predictors: (i) surprisal, frequency, and length, (ii) PMI, frequency, and length, and (iii) orthogonalized surprisal, frequency, and orthogonalized length. (An alternative perspective would be to use orthogonalized frequency, which assumes that surprisal is the more fundamental predictor. We provide such results, which are are consistent with the current conclusions, in § F.1.) We include spillover effects from the previous word and fit models over ten folds of cross-validation. We then measure the relative importance of predictors by averaging over the proportion of variance explained by the predictors across orderings in which they are added to the model (Lindeman et al., 1980; Kruskal, 1987), a technique known as LMG (Grömping, 2007). [^9] The intuition behind LMG is simple: Given a response variable ${\mathbf{y}}$ and two predictors ${\mathbf{x}_{1}}$ and ${\mathbf{x}_{2}}$, the proportion of the variance of ${\mathbf{y}}$ explained by ${\mathbf{x}_{1}}$ can be taken to either depend on the correlation between ${\mathbf{x}_{1}}$ and ${\mathbf{y}}$ (if ${\mathbf{x}_{1}}$ is added to the model before ${\mathbf{x}_{2}}$), or to depend on the partial correlation between ${\mathbf{x}_{1}}$ and ${\mathbf{y}}$ when controlling for ${\mathbf{x}_{2}}$ (if ${\mathbf{x}_{1}}$ is added to the model after ${\mathbf{x}_{2}}$). For a model with $p$ predictors, the LMG averages the explained variance over all $p!$ such orderings. One advantage of this technique, compared to previous methods (e.g., Delta Log Likelihood; Goodkind and Bicknell, 2018) is that LMG gives a better absolute sense of predictive power.
@@ Results.
The LMG values for the different predictors across these models are visualized in Fig. 1. Comparing the plots on the bottom row to the plots on the top row, we observe that the explained variance for orthogonalized surprisal is much lower than for surprisal. These results are consistent across languages. For most languages, when ranking the importance of the predictors, orthogonalization shifts the third most important predictor from context to the previous word’s frequency. Our results suggest that using surprisal therefore overestimates the importance that context has on reading time. [^10] Our results may help explain why surprisal estimates from larger LMs provide poorer fits to reading than those from medium-sized LMs (Oh and Schuler, 2023b); see App. G. We observe values for PMI that for most languages lie between those for surprisal and orthogonalized surprisal, indicating that the extent to which that PMI overestimates the context effect is smaller compared to surprisal. We observe that the mean $R^{2}$ values of the models across LMG orderings, i.e., the sums of the bars within each facet, range between 0.60.60.60.60.80.80.80.8, indicating that the linear models capture a fairly large proportion of the variance observed during the reading process.
@@ 5 Conclusion
Our approach takes one predictor to remain untouched (i.e., frequency), and modifies others to reflect effects that are disassociated from the first. As suggested above, it would thus be natural to ask what happens in the alternative setting, where surprisal remains untouched and frequency is projected onto the orthogonal complement of surprisal. We provide such an analysis in § F.1. It turns out that even when attributing the shared effect of frequency and surprisal on reading times to surprisal—which is the case when replacing frequency by an orthogonalized frequency predictor—the variance explained by the frequency predictor is still higher for most languages in comparison to the surprisal predictor. This gives further support to our conclusion that context appears to play a small role in reading time prediction.
Furthermore, our presentation of ideas and discussion largely ignores effects of word lengths. We find word lengths to explain the most variance in reading times in our surprisal and PMI models. In addition, after residualizing word length against frequency, we find length to be the second strongest predictor, with an explained variance ranging from around $10\%$$21\%$. One hypothesis is that readers may make multiple saccades within first passes of longer words, and the time it takes to plan and execute these saccades could be the underlying reason why orthogonalized word length remains explanatory even after residualization. Future work could control for this by adding in the number of saccades within a word as an additional predictor into models.
@@
Another limitation of this work is that, while we investigate several languages, these are still biased towards Indo-European languages. For example, we present results from one language only for Fino-Uralic, Semitic, Turkic, and Koreanic language families, but seven Indo-European languages. Expanding these results to even more languages would further broaden the impact of this work. In addition, we observe somewhat unique effects for Korean, where, in orthogonalized models, frequency accounts for a lower proportion of the variance, and length and context account for higher proportions, at least compared to other languages. One possible reason for this is the Korean script (Hangul), which combines features of both alphabetic and syllabic writing systems. Future work should conduct similar analyses on different Korean datasets to determine whether this trend is a property of Korean, or just our particular Korean language dataset.
@@ E.1 Dataset
We use the Multilingual Eye-movement Corpus (MECO; Siegelman et al., 2022), which consists of eye-tracking-based reading-time data across 13 different languages for 12 Wikipedia-style articles about various topics. The articles have been carefully constructed to contain the same content across languages. Word-level reading time is recorded for between $32{-}54$ participants per language, using several different reading variables. Of these, for our main experiments, we use gaze duration, which is the total fixation time on a word during its first pass (i.e., before the first time the gaze leaves the word). However, below we also show results for two other reading metrics: first fixation duration, which is the duration of the first fixation that lands on the word, and total fixation duration, which is the sum over all fixation durations of a word. In our experiments, we give a reading time of zero to words that were skipped on the first pass. We take the average over the by-subject reading times to obtain the response variables we model.
@@ Appendix F Additional Experiments
We complement the main experiments presented in § 4 with three additional empirical analyses. In § F.1, we take an alternative position to the one in the main text, using an orthogonalized frequency predictor. In § F.2, we exclude the length predictor and perform the same analysis as we did in the main text. In § F.3, we compare the predictive power of nonlinear models across different sets of predictors.
@@ F.1 Orthogonalized Frequency
Here, we provide an additional analysis in which we derive a new frequency predictor, swapping the two variables in Eq. 14. In this case, the shared effect of frequency and surprisal on reading time is attributed to surprisal, and the frequency effect represents the effect beyond what is already explained by surprisal. We present the results in Fig. 2. Comparing this analysis to Fig. 1 may be considered analogous to Shain’s (2019) study, which compares the independent effects of surprisal and frequency by adding them in as predictors on top of baseline models that contain the other. However, in contrast to Shain (2019), we find that frequency appears to be more important than contextual effects in explaining reading times.
@@ F.3 Psychometric Predictive Power
While the surprisal and PMI models are equivalent under a linear model, that relationship does not necessarily need to hold under a nonlinear one. Therefore, we compare the psychometric predictive power by fitting generalized additive models (GAMs). GAMs are a class of additive models that can learn non-linear relationships between predictor and response variables. All the terms in our GAMs are spline-based smooth terms, that include either a contextual predictor variable (i.e., surprisal, PMI, or orthogonalized surprisal), or frequency. We restrict our smooth terms to six basis functions, following the logic outlined in Hoover et al. (2023). GAMs are fit using the mgcv package in R. Two example calls are given below:
@@
Results are visualized in Fig. 4. We observe three big trends: First, we find that all predictors lead to positive $\Delta_{\mathrm{llh}}$, indicating that they are useful for predicting reading times. However, second, when looking at models with just one predictor variable, we observe that frequency alone leads to higher $\Delta_{\mathrm{llh}}$than any other single variable, and that surprisal and PMI tend to result in higher $\Delta_{\mathrm{llh}}$than orthogonalized surprisal. This is to be expected. We know from prior research that frequency plays a large role in explaining by-word processing effort, and because orthogonalized variables, by definition, are decorrelated with frequency, they are not expected to be strong predictors of reading times, alone.
@@
Figure 5: Psychometric Predictive Power by Language. Results are consistent those reported in Fig. 4.

### v4

@@ 2.2 Frequency as a Predictor of Reading Time
where the normalizing constant ${{Z}_{{q_{\mathrm{H}}}}}$ is necessarily finite for language models of finite expected length. [^3] As a simple counterexample, consider an LM ${{p_{\mathrm{H}}}}$ with ${\Sigma}=\{a\}$ and ${{p_{\mathrm{H}}}}(a^{n})=\frac{1}{\pi^{2}/6}\cdot\frac{1}{n^{2}}$ for $n\in\mathbb{Z}_{>0}$, i.e., where the probabilities are globally normalized by $\frac{\pi^{2}}{6}$, the solution to the Basel problem. The expected count of $a$ would depend on $\sum_{n=1}^{\infty}\frac{1}{n^{2}}\cdot n$, which is divergent. Thus, ${{Z}_{{q_{\mathrm{H}}}}}=\infty$. Then, given an LM ${{p_{\mathrm{H}}}}$ with a unigram LM ${{q_{\mathrm{H}}}}$, the unigram surprisal is given by
@@ 4 Variance Explained by Context
We fit three linear models using ordinary least squares, with predictors: (i) surprisal, frequency, and length, (ii) PMI, frequency, and length, and (iii) orthogonalized surprisal, frequency, and orthogonalized length. (An alternative perspective would be to use orthogonalized frequency, which assumes that surprisal is the more fundamental predictor. We provide such results, which are are consistent with the current conclusions, in § F.1.) We include spillover effects from the previous word and fit models over ten folds of cross-validation. We then measure the relative importance of predictors by averaging over the proportion of variance explained by the predictors across orderings in which they are added to the model (Lindeman et al., 1980; Kruskal, 1987), a technique known as LMG (Grömping, 2007). [^9] The intuition behind LMG is simple: Given a response variable ${\mathbf{y}}$ and two predictors ${\mathbf{x}_{1}}$ and ${\mathbf{x}_{2}}$, the proportion of the variance of ${\mathbf{y}}$ explained by ${\mathbf{x}_{1}}$ can be taken to either depend on the correlation between ${\mathbf{x}_{1}}$ and ${\mathbf{y}}$ (if ${\mathbf{x}_{1}}$ is added to the model before ${\mathbf{x}_{2}}$), or to depend on the partial correlation between ${\mathbf{x}_{1}}$ and ${\mathbf{y}}$ when controlling for ${\mathbf{x}_{2}}$ (if ${\mathbf{x}_{1}}$ is added to the model after ${\mathbf{x}_{2}}$). For a model with $p$ predictors, the LMG averages the explained variance over all $p!$ such orderings. One advantage of this technique, compared to previous methods (e.g., delta log-likelihood; Goodkind and Bicknell, 2018) is that LMG gives a more interpretable absolute score for predictive power.
@@ Results.
The LMG values for the different predictors across these models are visualized in Fig. 1. Comparing the plots on the bottom row to the plots on the top row, we observe that the explained variance for orthogonalized surprisal is much lower than for standard contextual surprisal. These results are consistent across languages and suggest that using surprisal overestimates the importance that context has on reading time. We observe values for PMI that for most languages lie between those for surprisal and orthogonalized surprisal, indicating that the extent to which that PMI overestimates the context effect is smaller compared to surprisal. Moreover, we note that the context-based predictors are considerably less predictive than frequency and (orthogonalized) length. [^10] Our results may help explain why surprisal estimates from larger LMs provide poorer fits to reading than those from medium-sized LMs (Oh and Schuler, 2023b); see App. G. This holds for all languages except Korean. Excluding Korean, we observe that the average $R^{2}$ values of the models across LMG orderings, i.e., the sums of the bars within each facet, range between 0.250.250.250.250.450.450.450.45, indicating that the linear models capture a moderate proportion of the variance in gaze duration.
@@ 5 Conclusion
Our approach takes one predictor to remain untouched (i.e., frequency), and modifies others to reflect effects that are disassociated from the first. As suggested above, it would thus be natural to ask what happens in the alternative setting, where surprisal remains untouched and frequency is projected onto the orthogonal complement of surprisal. We provide such an analysis in § F.1. It turns out that even when attributing the shared effect of frequency and surprisal on reading times to surprisal—which is the case when replacing frequency by an orthogonalized frequency predictor—the variance explained by the frequency predictor is still higher for most languages in comparison to the surprisal predictor. This gives further support to our conclusion that context appears to play a smaller role in reading time prediction.
Furthermore, our presentation of ideas and discussion largely ignores effects of word lengths. We find word lengths to explain the most variance in reading times in our surprisal and PMI models. In addition, after residualizing word length against frequency, we find length to be the second strongest predictor, with an explained variance ranging from around $7\%$$12\%$. One hypothesis is that readers may make multiple saccades within first passes of longer words, and the time it takes to plan and execute these saccades could be the underlying reason why orthogonalized word length remains explanatory even after residualization. Future work could control for this by adding in the number of saccades within a word as an additional predictor into models.
@@
Another limitation of this work is that, while we investigate several languages, these are still biased towards Indo-European languages. For example, we present results from one language only for Fino-Uralic, Semitic, Turkic, and Koreanic language families, but seven Indo-European languages. Expanding these results to even more languages would further broaden the impact of this work. In addition, we observe somewhat unique effects for Korean, where frequency accounts for a lower proportion of the variance, and length and context account for higher proportions. One possible reason for this is the Korean script (Hangul), which combines features of both alphabetic and syllabic writing systems. Future work should conduct similar analyses on different Korean datasets to determine whether this trend is a property of Korean, or just our particular Korean language dataset.
@@ E.1 Dataset
We use the Multilingual Eye-movement Corpus (MECO; Siegelman et al., 2022), which consists of eye-tracking-based reading-time data across 13 different languages for 12 Wikipedia-style articles about various topics. The articles have been carefully constructed to contain the same content across languages. Word-level reading time is recorded for between $32{-}54$ participants per language, using several different reading variables. Of these, for our main experiments, we use gaze duration, which is the total fixation time on a word during its first pass (i.e., before the first time the gaze leaves the word). However, in Fig. 5 below we also show results for two other reading metrics: first fixation duration, which is the duration of the first fixation that lands on the word, and total fixation duration, which is the sum over all fixation durations of a word. In our experiments, we average the reading times for each word across all participants, excluding the participants that skipped (i.e., did not fixate on) the word. Thus, we avoid modeling the variance stemming from whether a word is skipped.
@@ Appendix F Additional Experiments
We complement the main experiments presented in § 4 with three additional empirical analyses. In § F.1, we take an alternative position to the one in the main text, using an orthogonalized frequency predictor. In § F.2, we exclude the length predictor and perform the same analysis as we did in the main text. In § F.3, we compare the predictive power of nonlinear models across different sets of predictors, including analyses on first fixation duration and total fixation duration.
@@ F.1 Orthogonalized Frequency
Here, we provide an additional analysis in which we derive a new frequency predictor, swapping the two variables in Eq. 14. In this case, the shared effect of frequency and surprisal on reading time is attributed to surprisal, and the frequency effect represents the effect beyond what is already explained by surprisal. We present the results in Fig. 2. Comparing this analysis to Fig. 1 may be considered analogous to Shain’s (2019) study, which compares the independent effects of surprisal and frequency by adding them in as predictors on top of baseline models that contain the other. However, in contrast to Shain (2019), we find that frequency appears to be more important than contextual effects in explaining reading times; for most languages, frequency remains the more important predictor even after orthogonalization.
@@ F.3 Psychometric Predictive Power
While the surprisal and PMI models are equivalent under a linear model, that relationship does not necessarily need to hold under a nonlinear one. Therefore, we compare the psychometric predictive power by fitting generalized additive models (GAMs). GAMs are a class of additive models that can learn non-linear relationships between predictor and response variables. All the terms in our GAMs are spline-based smooth terms, that include either a contextual predictor variable (i.e., surprisal, PMI, or orthogonalized surprisal), or frequency. We restrict our smooth terms to six basis functions, following the logic outlined by Hoover et al. (2023). GAMs are fit using the mgcv package in R. Two example calls are given below:
@@
Results for gaze duration are visualized in Fig. 4. We observe three big trends: First, we find that all predictors lead to positive $\Delta_{\mathrm{llh}}$, except for the case of Korean, indicating that they are useful for predicting reading times. However, second, when looking at models with just one predictor variable, we observe that frequency alone leads to higher $\Delta_{\mathrm{llh}}$than surprisal and PMI do, and that surprisal and PMI tend to result in higher $\Delta_{\mathrm{llh}}$than orthogonalized surprisal. This is to be expected. We know from prior research that frequency plays a large role in explaining by-word processing effort, and because orthogonalized variables, by definition, are decorrelated with frequency, they are not expected to be strong predictors of reading times, alone.
@@
Figure 5: $\Delta_{\mathrm{llh}}$of GAM models over three different aggregated measurements of reading time. Error bars are 95% confidence intervals across folds of data.